# Homogenized C. elegans Neural Activity and Connectivity Data

### v3

@@ 2.3.2 Data Extraction and Preprocessing
    Supplementing Attributes: Some desirable attributes about the connectome (e.g., neuron positions, neurotransmitters, cell types, etc.) may not available from the source datasets. We obtain these from published research (e.g., neuron positions from [16], neurotransmitters and cell types from [27]) to supplement our graph structures.
@@ 2.3.3 Graph Tensor Format
To facilitate downstream analyses and modeling, we standardized the connectome data into a graph tensor format compatible with graph neural network (GNN) libraries. Each dataset is transformed into a torch_geometric [21] Data object with the following components:
@@ 2.3.4 Consensus Connectome
To address discrepancies in counts and weights of connections between datasets, we propose a consensus connectome that aggregates data across sources using majority voting. For each edge $e_{ij}$, the consensus weights are:
$c_{\text{consensus}}=\max\{c_{1},c_{2},\dots,c_{n}\},\quad g_{\text{consensus}}=\max\{g_{1},g_{2},\dots,g_{n}\},$
where $c_{i},g_{i}$ are the chemical and gap junction weights from $i$-th connectome source.
In order to quantify the uncertainty of this connectome, we also compute a weighted standard deviation $v$ of each edge across sources and include this as an additional attribute $\mathbf{a}=[c,g,...,v]$:
$v_{\text{consensus}}=\lambda\cdot\operatorname{std}\{c_{1},c_{2},\dots,c_{n}\}+(1-\lambda)\cdot\operatorname{std}\{g_{1},g_{2},\dots,g_{n}\}$
where $\lambda$ is a tuneable hyperparameter determining the relative importance of the chemical versus electrical synapse weights. We set $\lambda=0.9$ reflecting our knowledge that annotations of gap junctions are less certain than those of chemical synapses and that gap junctions are more sparse than chemical synapses in the connectome [3].
Having a single consensus connectome could used to model the average C. elegans nervous system, which has a stereotyped neurodevelopmental trajectory [35], or to quantify the inter-animal variability in network weights, which may reflect environmental influences. Figure 5 visualizes the consensus connectome of the 22 amphid chemosensory neurons [11].
@@ 3 Discussion
In this paper we present a normalized and standardized datasets for C. elegans neural activity and connectomes. We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code is composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
The homogenized neural and connectome datasets provide a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. While our datasets provide a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using these datasets. There are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models. We address most most of these in the next few paragraphs.
@@
While the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and across the lifespan of a single individual [36]. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [14]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity measurements. This creates both an opportunity and a limitation. The opportunity arises from the fact that any robust relationship between structure and function to be discovered from this data must generalize across different age and phenotype-matched animals. The limitation is that this may preclude the development of individual-specific models which are desirable for understanding learning and memory, for example. While having the C. elegans connectome [35, 3] provides a comprehensive structural map of synaptic connections, understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons [14] due to neuromodulation by way of neuropeptides which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [1, 31].
We opted not to apply smoothing to the neural activity traces for two principal reasons. First, it is likely that the original data creators have already undertaken a degree of preprocessing on their raw microscopy or imaging data, which often includes smoothing techniques to enhance signal quality. Introducing additional smoothing could risk over-processing the data, thereby obscuring critical neural dynamics. Second, the calcium fluorescence signals utilized in our analysis are inherently low-pass filtered representations of the underlying neural activity, a characteristic of the imaging modality itself. This filtering effectively diminishes high-frequency components, which are vital for capturing rapid neural events. By refraining from further smoothing, we aim to preserve the integrity and richness of the original data, allowing for a more nuanced understanding of the biological phenomena under investigation.
@@
One important clarification about our work needs to made regarding our use of $300$ for the number of neurons in C. elegans, whereas readers may be more familiar the number $302$ from prior literature. We drop the bilateral pair of canal-associated neurons (CAN) from our datasets because they completely lack chemical synapses and are now widely classified as ‘end-organs’ rather than neurons [17, 3]. The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections. Moreover, the used of calcium measurements might bias neural activity models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions. Researchers should be mindful of these constraints when developing models and consider integrating additional multimodal datasets to build more comprehensive models [8].
Despite these limitations, the C. elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics. For instance, graph neural networks (GNNs) [34] are well-suited to handle the structural connectivity data, while time-series models like structured state space models [25] can be applied to the neural activity data to capture temporal dependencies. These models could inform the development of foundation models that generalize to larger and more complex systems. Moreover, the datasets’ open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
@@ 4 Conclusion
We present a standardization protocol implemented in a code repository for the preprocessing and integration of 11 C. elegans calcium fluorescence datasets and 3 primary connectome sources [35, 3, 36]. The code repository contains preprocessing classes that load and preprocess the different source datasets, handling both calcium fluorescence and connectome data. Each dataset has a class implementation that inherits from a respective parent class, which has been adapted to accommodate varying raw data formats. Once loaded, the calcium fluorescence data is further masked, normalized, resampled, and smoothed.

### v4

@@ 2.3.2 Data Extraction and Preprocessing
    Supplementing Attributes: Some desirable attributes about the connectome (e.g., neuron positions, neurotransmitters, cell types, etc.) may not be available from the source datasets. We obtain these from published research (e.g., neuron positions from [17], neurotransmitters and cell types from [28]) to supplement our graph structures.
@@ 2.3.3 Graph Tensor Format
To facilitate downstream analyses and modeling, we standardized the connectome data into a graph tensor format compatible with the PyTorch Geometric graph neural network (GNN) libraries. Each dataset is transformed into a torch_geometric [22] Data object with the following components:
@@ 2.3.4 Consensus Connectome
To address discrepancies in connection counts and weights across datasets, we construct a consensus connectome that aggregates synaptic data from multiple sources using a directional averaging strategy. Unlike the dataset-specific files used during preprocessing — which store information in graph network format with edge indices and attributes — the consensus connectome is a flat .csv table, not a graph. This tabular format is universal, self-explanatory, and easily readable by both humans and software tools. Each row corresponds to a directed pair of neurons, with clearly defined anatomical, connection, and metadata fields. This structure is especially suited for accessible downstream analysis, visualization, and integration, making it the preferred format for summarizing the overall state of connectivity across datasets.
For each ordered pair of neurons $(i,j)$, we compute the consensus weights as follows:
  • Gap Junctions are symmetric. Their consensus weight $g_{\text{consensus}}$ is computed as the mean of all non-zero gap weights observed in either direction across datasets:
    $g_{\text{consensus}}=\operatorname{mean}\{g_{ij}^{(k)}\}$
    where $g_{ij}^{(k)}$ is the gap weight from neuron $i$ to $j$ in dataset $k$. By definition, $g_{ij}^{(k)}=g_{ji}^{(k)}$.
  • Chemical Synapses are directional. Their consensus weight $c_{\text{consensus}}$ is the mean of all non-zero chemical weights from neuron $i$ to $j$, excluding values from the functional connectivity dataset:
    $c_{\text{consensus}}=\operatorname{mean}\{c_{ij}^{(k)}\}_{k\neq\text{funconn}}.$
  • Functional Connectivity is reported directly from the dedicated funconn dataset, if available. These values are not averaged and apply only in the directed case $i\rightarrow j$.
To quantify the uncertainty in gap and chemical connection estimates, we compute a weighted standard deviation $v$ that reflects both gap and chemical synapse variability:
$v_{\text{consensus}}=\lambda\cdot\operatorname{std}(\{c_{ij}^{(k)}\}_{k\neq\text{funconn}})+(1-\lambda)\cdot\operatorname{std}(\{g_{ij}^{(k)}\}_{k}),$
where $\lambda=0.9$ is a subjective value chosen to weight the uncertainty from chemical synapses more due to their relative abundance compared to gap junctions.
This consensus connectome includes all neuron pairs — including self-loops — and supports multiple output formats (e.g., trimmed vs. full, with or without functional weights). It provides a representative model of the stereotyped C. elegans nervous system [37], while capturing cross-dataset variability that may reflect biological and technical noise. Figure 5 visualizes the consensus connectivity among the 22 amphid chemosensory neurons [12].
@@ 3 Discussion
In this paper we present standardized datasets for C. elegans neural activity and connectomes. We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code is composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
The homogenized neural and connectome datasets provide a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. While our datasets provide a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using these datasets. There are several important limitations and methodological choices that must be considered when interpreting the data and applying it in bio-computational models. We address most of these next.
@@
While the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and across the lifespan of a single individual [38]. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [15]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity measurements. This creates both an opportunity and a limitation. The opportunity arises from the fact that any robust relationship between structure and function to be discovered from this data must generalize across different age and phenotype-matched animals. The limitation is that this may preclude the development of individual-specific models which are desirable for understanding learning and memory, for example. While having the C. elegans connectome [37, 4] provides a comprehensive structural map of synaptic connections, understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons [15] due to neuromodulation by way of neuropeptides which can bind to nearby and distant neuron GPCR sites triggering secondary messenger pathways [2, 33].
We opted not to apply smoothing to the neural activity traces for two principal reasons. First, it is likely that the original data creators have already performed a significant amount preprocessing on the signals extracted from their raw microscopy images, which often includes smoothing techniques to enhance signal quality. Introducing additional smoothing could risk over-processing the data, obscuring critical neural dynamics. Second, the calcium fluorescence signals utilized in our analysis are inherently low-pass filtered representations of the underlying neural activity, a characteristic of the imaging modality itself. This filtering effectively diminishes high-frequency components, which are vital for capturing rapid neural events. By refraining from further smoothing, we aim to preserve the integrity and richness of the original data, allowing for a more nuanced understanding of the biological phenomena under investigation.
@@
One important clarification about our work needs to be made regarding our use of $300$ for the number of neurons in C. elegans, whereas readers may be more familiar the number $302$ from prior literature. We drop the bilateral pair of canal-associated neurons (CAN) from our datasets because they completely lack chemical synapses and are now widely classified as ‘end-organs’ rather than neurons [18, 33]. The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections. Moreover, the use of calcium measurements might bias neural activity models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions. Researchers should be mindful of these constraints when developing models and consider integrating additional multimodal datasets to build more comprehensive models [9].
Despite these limitations, the C. elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics. For instance, graph neural networks (GNNs) [36] are well-suited to handle the structural connectivity data, while time-series models like structured state space models (SSMs) [26] can be applied to the neural activity data to capture temporal dependencies. These models could inform the development of foundation models that generalize to larger and more complex systems. Moreover, the datasets’ open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
@@ 4 Conclusion
We present a standardization protocol implemented in a code repository for the preprocessing and integration of 12 C. elegans calcium fluorescence datasets and 4 primary connectome publications [37, 4, 38, 15]. The code repository contains preprocessing classes that load and preprocess the different source datasets, handling both calcium fluorescence and connectome data. Each dataset has a class implementation that inherits from a respective parent class, which has been adapted to accommodate varying raw data formats. Once loaded, the calcium fluorescence data is further masked, normalized, resampled, and smoothed.